OpenStudy (anonymous):

show me how to solve 4y+15=6y-11

OpenStudy (anonymous):

\[4y+15=6y-11\]

OpenStudy (anonymous):

\[15=6y-4y-11\]

OpenStudy (anonymous):

\[15+11=2y\]

OpenStudy (anonymous):

\[26=2y\]

OpenStudy (anonymous):

\[26/2=y\]

OpenStudy (anonymous):

4y+15=6Y-11 -15 -15 4y=6y =26

OpenStudy (anonymous):

\[y=13\]

OpenStudy (anonymous):

4y=6y + 26

OpenStudy (anonymous):

4y=6y+26 -6y -6y -2y= 26 divide by 2 on both sides and you get -13

OpenStudy (anonymous):

I dont think you're correct @santosrubi_5

OpenStudy (anonymous):

where did i mess up?

OpenStudy (anonymous):

Umm..@santosrubi_5 you wrote this: 4y+15=6Y-11 -15 -15 4y=6y+26

OpenStudy (anonymous):

If you're solving that way, it should be: 4y+15=6Y-11 -15 -15 4y=6y-26

OpenStudy (anonymous):

Yes, youre right:)

OpenStudy (anonymous):

Ok, so @santosrubi_5 you agree that y=13?

OpenStudy (anonymous):

yes i do